OpenStudy (anonymous):

(36^t6 - 27t^5 -9t ^2) / (-9t^2) 9t^ (4t^4 - 3t^3 -1)/-9t^ 9t^2 cancel out and the answer is 4t^4 - 3t^3 -1 Just want to know if I am correct

OpenStudy (anonymous):

\[\frac{(36t^6 - 27t^5 -9t ^2)}{(-9t^2)} =\frac{9t ^2(4t^4 - 3t^3 -1)}{(-9t^2)}=-(4t^4 - 3t^3 -1)\] you had just missed the minus sign before your ans

OpenStudy (anonymous):

Thank you so very much =)